OpenStudy (anonymous):

Which of the following choices is the simple formula for the nth term of the following arithmetic sequence? 5, 1, -3, -7, -11, ... A. -4n + 1 B. 5n + 4 C. -4n + 9 D. 4n + 9 E. 4n + 1

OpenStudy (er.mohd.amir):

a=5 d=-11-(-7)=-4 n th term=a+(n-1)d solve

OpenStudy (michele_laino):

I think that the cnstant of such sequence is \(d=-4\), so the n-th term is given by the subsequent formula: \[\begin{gathered} {a_n} = {a_1} + \left( {n - 1} \right)\left( { - 4} \right) = \hfill \\ \hfill \\ = 5 + \left( {n - 1} \right)\left( { - 4} \right) = ...? \hfill \\ \end{gathered} \] namely, the same answer of @Er.Mohd.AMIR

OpenStudy (hope210):

I agree with @Er.Mohd.AMIR and @Michele_Laino

OpenStudy (anonymous):

okay so it would be 4n+9

OpenStudy (er.mohd.amir):

5-4n+4 solve

OpenStudy (anonymous):

it would be -4n+1

OpenStudy (er.mohd.amir):

-4n+9

OpenStudy (anonymous):

i mean 9 right

OpenStudy (anonymous):

it hoguht the 5 was a negitive at first lol

OpenStudy (er.mohd.amir):

ok

OpenStudy (anonymous):

ty